rxcklesskaisher:

Which of the values in the set {2, 3, 4, 5} is a solution to the equation 2x + 4 = 10? A) 2 B) 3 C) 4 D) 5

mdobbs6856:

2x+4=10 -4 -4 2x=6 x=3

rxcklesskaisher:

the correct answer is 3?

mdobbs6856:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher the correct answer is 3? \(\color{#0cbb34}{\text{End of Quote}}\) Yes!